OpenStudy (anonymous):

1. Which of the following is a rational number?

OpenStudy (anonymous):

remeber me lolfun i just changed to @lovepurle

OpenStudy (anonymous):

\sqrt{102} or \sqrt{29 } \sqrt{121 } \]

OpenStudy (anonymous):

Uh?

OpenStudy (anonymous):